# Hints of Prompt: Enhancing Visual Representation for Multimodal LLMs in Autonomous Driving

### v1

@@ Abstract
In light of the dynamic nature of autonomous driving environments and stringent safety requirements, general MLLMs combined with CLIP alone often struggle to represent driving-specific scenarios accurately, particularly in complex interactions and long-tail cases. To address this, we propose the Hints of Prompt (HoP) framework, which introduces three key enhancements: Affinity hint to emphasize instance-level structure by strengthening token-wise connections, Semantic hint to incorporate high-level information relevant to driving-specific cases, such as complex interactions among vehicles and traffic signs, and Question hint to align visual features with the query context, focusing on question-relevant regions. These hints are fused through a Hint Fusion module, enriching visual representations and enhancing multimodal reasoning for autonomous driving VQA tasks. Extensive experiments confirm the effectiveness of the HoP framework, showing it significantly outperforms previous state-of-the-art methods across all key metrics.
00footnotetext: Work conducted as an intern at DeepRoute.AI. †Corresponding author.
@@ 1 Introduction
In recent years, there have been growing trends towards formalizing autonomous driving within end-to-end learning-based frameworks [18]. Leveraging extensive human driving data, these approaches demonstrate human-like behavior and generalization capabilities, enabling them to handle intricate scenarios more effectively. However, the end-to-end pipeline often functions as a black box, making it challenging to interpret or explain the neural network-driven planning processes [7]. This opacity raises significant ethical and legal concerns regarding accountability and trustworthiness in safety-critical applications.
Interpretability has emerged as a crucial research focus in autonomous driving to address these concerns. Visual Question Answering (VQA) tasks have gained considerable attention in this context [39], as they allow systems to articulate their observations and actions through natural language interactions. By bridging the gap between complex decision-making processes and user comprehension, VQA enhances system transparency and fosters user trust.
@@
The application of Multimodal Large Language Models (MLLMs) [4, 46, 42, 8, 26, 14] has significantly advanced VQA tasks by integrating visual encoders with Large Language Models (LLMs) [12, 43, 10] for deeper multimodal understanding. Typically, MLLMs comprise a visual encoder, an adapter, and an LLM. For VQA, the visual encoder extracts image features, aligned by the adapter with text tokens. The LLM tokenizes the question and integrates the aligned visual tokens to generate an answer.
@@
For VQA in autonomous driving, the dynamic nature of driving environments and stringent safety requirements pose challenges for general MLLMs, particularly in representing hierarchical visual features crucial for complex interactions and long-tail cases. Although CLIP-based models [37, 50] have dominated visual encoders, as shown in Fig. 2, CLIP visual tokens overlook token-wise affinity, which is essential for preserving instance-level structure and establishing basic objectness understanding in driving scenarios. In contrast, DINOv2 [35] better maintains this structure. Additionally, general visual encoders underrepresent domain-specific semantics, making them prone to missing small yet crucial elements for driving, such as distant vehicles, pedestrians, and traffic signs—features vital for reasoning in complex interactions, like pedestrians unexpectedly crossing the street. Lastly, current MLLMs struggle to adjust visual representations based on specific questions, as visual and text tokens are processed separately. This separation limits the model’s ability to focus on query-relevant image regions, reducing the effectiveness of context-aware responses in autonomous driving.
@@
Based on these observations, we introduce stratified tokens as hints to enhance visual representation. Specifically, we propose three key enhancements: Affinity hint to capture instance-level structure and strengthen token-wise connections, Semantic hint to incorporate high-level semantic information for domain-specific elements, and Question hint to correlate visual features according to the query context, ensuring focus on question-relevant regions. We conduct comprehensive studies on designing effective and efficient hints fusion strategies for combining question-specific and hierarchical semantic cues tailored to autonomous driving, thereby enhancing multimodal reasoning in these scenarios. Moreover, we propose an efficient version of HoP that maintains competitive performance, achieved by extracting Affinity and Semantic hints directly from two lightweight heads distilled and fine-tuned from the CLIP backbone. Fig. 1 illustrates our proposed method, referred to as Hints of Prompt (HoP) framework throughout the paper.
Extensive experiments demonstrate significant improvements for VQA. On the LingoQA [34] benchmark, as shown in Sec. 3.3, HoP achieves a Lingo-Judge score of 67.8, establishing a new state-of-the-art. Ablation studies confirm that the three types of hints are complementary, collectively enhancing representation capability. Additionally, we explored various forms of Semantic and Affinity hints from different sources. Results indicate that as long as hint tokens represent token-wise affinity, performance improves, even when using a similarity matrix as Affinity hint tokens.
To summarize, we would like to highlight our contributions as follows:
@@
  • We identify key limitations in current MLLMs combined with CLIP alone for VQA tasks in autonomous driving.
  • We introduce the Hints of Prompt (HoP) framework, which incorporates Affinity, Semantic, and Question hint tokens through a fusion module, significantly enhancing visual representation, enriching the LLM’s understanding of complex driving scenes, and boosting overall performance.
@@
  • Extensive experiments demonstrate significant improvements of our proposed methods for VQA, establishing new state-of-the-art results on LingoQA [34], DRAMA [32], and BDD-X [21].
@@ 2 Related Work
Figure 3: A VQA sample case visualizes the impact of the three proposed hint types compared to the baseline.
@@
In autonomous driving, MLLMs address critical challenges in perception, planning, control, question answering (QA), and generative world models. For example, GPT-Driver [33] reframes motion planning as a language modeling task, enhancing accuracy, while DriveGPT4 [48] combines vision and language inputs to predict control signals with real-time explanations, improving transparency.
In perception, MLLMs advance sensor data fusion from LiDAR and cameras; models like Talk2BEV [11] use Bird’s Eye View (BEV) features for complex visual tasks, while LMDrive [38] and DriveVLM [41] employ trajectory-based reasoning to refine scene understanding and decision-making. In QA, DriveLM [40] uses graph-based reasoning to enhance real-time planning. For world models, generative approaches like DriveDreamer [47] and GAIA-1 [19] generate realistic driving scenarios for data augmentation and training. These advancements highlight MLLMs’ potential to enhance multimodal understanding and optimize autonomous driving systems across diverse applications.
Enhancing Visual Representations in MLLMs. Effective alignment of visual representations with LLM capabilities is essential for improving overall performance. Some approaches focus on enhancing visual encoders directly. For instance, InternVL [8] introduces a large-scale visual encoder to improve feature granularity, while DeepSeek-VL [30] and Cambrian-1 [42] combine multiple encoders like SigLIP [50] and SAM [22] to leverage complementary strengths. To improve efficiency, BLIP-2 [23] employs a lightweight querying transformer to bridge vision and language models.
QA-ViT [17] incorporates textual information into visual features to help MLLMs focus on relevant image regions, enhancing interpretability while maintaining efficiency. Similarly, we utilize the same LLM text embeddings as Question hints, but our fusion strategy is more easily integrated with various visual backbones. Hint-AD [15] introduces a holistic alignment between natural language and intermediate perception and planning tokens, improving interpretability in complex scenarios. In contrast, we further introduce instance-level and question-level hint tokens with a lightweight fusion module, enabling the MLLM visual encoder to dynamically extract relevant hints for a better understanding of the current driving scenario.
@@ 3.1 Overview
As shown in Fig. 1, the Hints of Prompt (HoP) framework leverages three distinct token types—Affinity, Semantic, and Question hints—each representing a different level of information. These hints progress from instance-level, token-wise connections in the Affinity hint, to high-level semantic information in the Semantic hint, and finally to question-specific context in the Question hint, gradually refining and aligning the visual representation with the LLM. A simple Hint Fusion module integrates these hints into a cohesive visual representation, which is then processed by an adapter and passed to the LLM for answer generation. By enabling the visual encoder to dynamically incorporate hints across different information levels, our approach enhances context-aware responses in VQA for autonomous driving scenarios.
@@ 3.2 Hint Tokens
Semantic hint After implementing instance-level enhancement, providing a semantic category specific to the driving domain can further focus visual representations on key instances, such as the cyclist shown in Fig. 3. The Semantic hint aims to incorporate high-level, domain-specific semantic information, such as vehicles, traffic signs, and pedestrians, which is crucial for understanding complex interactions. To achieve this, we employ sparse queries from DETR-like [6] detection or segmentation pipelines as Semantic hints. Specifically, GroundDINO [28] or Mask2Former [9] will be utilized for Semantic hint extraction. Experimental results (Tab. 5) show that, benefiting from Mask2Former’s adaptation to driving scenarios, it aligns more effectively with the task-specific requirements of VQA in autonomous driving.
@@
As shown in Fig. 3, incorporating the three proposed hints significantly improves the model’s response over the baseline. The baseline incorrectly suggests overtaking is possible, overlooking oncoming traffic. With the added hints, the model gains instance-level structure and focuses more accurately on key elements, such as the cyclist and oncoming vehicles. This added context enhances the model’s spatial reasoning, resulting in a more precise and relevant response in complex driving scenarios. Detailed quantitative analysis is available in Sec. 4.
@@ 3.3 Hints Fusion
Concatenation As a baseline strategy, hint tokens are directly concatenated with CLIP tokens, as illustrated in Fig. 4(a):
@@
{booktabs}
colsep = 11pt, colspec = lccccc, row8-9 = bg=gray!15, cell12-6 = c, Method LLM Lingo-Judge $\uparrow$ BLEU-4 $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$
GPT-4V [49] (zero-shot) - 59.6 6.30 12.4 42.8
LingoQA Baseline [34] Vicuna-v1.5-7B 60.6 15.0 18.6 65.6
QA-ViT [17]† Vicuna-v1.5-7B 63.6 14.6 19.0 65.9
VTS [31] InternLM2-7B 64.2 14.5 20.5 56.9
LLaVA-v1.5† Vicuna-v1.5-7B 63.2 14.1 19.3 63.7
LLaVA-v1.5 (DINOv2)† Vicuna-v1.5-7B 62.8 14.7 19.3 68.3
Efficient HoP Vicuna-v1.5-7B 66.8 15.2 20.0 66.2
HoP Vicuna-v1.5-7B 67.8 15.8 20.3 70.9
@@ 4 Experiments
Table 2: Performance comparison on the DRAMA dataset. “†” indicates model that are fine-tuned on the DRAMA dataset. HoP demonstrates enhanced performance over other methods across most metrics, with LLaVA-v1.5 as the baseline.
{booktabs}
[…]
@@ 4.1 Implementation Details
The whole HoP training is conducted for one epoch using the AdamW optimizer with a cosine learning rate schedule, a warm-up ratio of 0.03, and a base learning rate of 2e-5. All models are trained on 32 NVIDIA A100 GPUs with a batch size of 4 per GPU.
Figure 5: Impact of each hint type within the HoP framework. Abbreviations: AH (Affinity hint), SH (Semantic hint), QH (Question hint). “w/ cls” indicates the addition of class information to the Semantic hint, and “ALL” denotes the combination of all three hint types.
@@ 4.2 Datasets and Benchmarks
LingoQA [34] is a benchmark dataset for video question answering in autonomous driving, with 28K unique driving scenarios and 419K question-answer (QA) pairs across tasks like object recognition, scene description, driving reasoning, and action justification. The dataset consists of two subsets: Action and Scenery, plus a high-quality Evaluation subset for benchmarking. The main metric, Lingo-Judge, is an evaluation model aligned with human preferences, while BLEU [36], METEOR [5], and CIDEr [45] provide lexical similarity assessments.
@@
BDD-X [21] is a dataset developed for autonomous driving tasks, focusing on video captioning and rationales, but it lacks native VQA pairs. To address this, we adopt the VQA pairs and evaluation protocol provided by DriveGPT4 [48], which include 26K training samples and 1.8K testing samples. Metrics such as BLEU, CIDEr, and ROUGE are used to ensure consistency and comparability of results.
4.3 Comparison with Existing Works
Our results on the LingoQA [34], DRAMA [32], and BDD-X [21] datasets demonstrate the effectiveness of the Hints of Prompt (HoP) method in enhancing VQA for autonomous driving. As shown in Sec. 3.3, HoP significantly outperforms baseline models on LingoQA, achieving a top Lingo-Judge score of 67.8 and leading BLEU-4, METEOR, and CIDEr scores, reflecting strong semantic accuracy and alignment with ground truth answers.
As shown in Sec. 4, HoP surpasses other methods on the DRAMA dataset across key metrics, including BLEU-1, BLEU-4, METEOR, ROUGE, and SPICE, highlighting its capability to generate detailed and contextually accurate answers in urban traffic scenes.
In Sec. 4, HoP’s performance on BDD-X, evaluated using the DriveGPT4 protocol, consistently outperforms other methods across all difficulty levels (Easy, Medium, Hard, and All), achieving the highest CIDEr scores and robust results even in challenging driving scenarios.
Visualization examples can be found in Fig. 7.
@@ 4.3 Comparison with Existing Works
Table 5: Impact of different Semantic hint types on performance. Abbreviations: SH (Semantic hint), LJ (Lingo-Judge), B4 (BLEU-4).
@@
Table 6: Impact of different Question hint types on performance. “CLIP” refers to hint tokens derived from the CLIP text encoder, while “LLM” refers to tokens from the LLM input embedding layer. Abbreviations: QH (Question Hint), LJ (Lingo-Judge), B4 (BLEU-4).

### v2

@@ Abstract
In light of the dynamic nature of autonomous driving environments and stringent safety requirements, general MLLMs combined with CLIP alone often struggle to accurately represent driving-specific scenarios, particularly in complex interactions and long-tail cases. To address this, we propose the Hints of Prompt (HoP) framework, which introduces three key enhancements: Affinity hint to emphasize instance-level structure by strengthening token-wise connections, Semantic hint to incorporate high-level information relevant to driving-specific cases, such as complex interactions among vehicles and traffic signs, and Question hint to align visual features with the query context, focusing on question-relevant regions. These hints are fused through a Hint Fusion module, enriching visual representations by capturing driving-related representations with limited domain data, ensuring faster adaptation to driving scenarios. Extensive experiments confirm the effectiveness of the HoP framework, showing that it significantly outperforms previous state-of-the-art methods in all key metrics.
@@ 1 Introduction
In recent years, there have been growing trends towards formalizing autonomous driving within end-to-end learning-based frameworks [18]. These approaches use extensive human driving data to demonstrate human-like behavior and generalization capabilities, enabling them to handle complex scenarios more effectively. However, the end-to-end pipeline often functions as a black box, making it challenging to interpret or explain the planning processes driven by neural networks [7]. This opacity raises significant ethical and legal concerns regarding accountability and trustworthiness in safety-critical applications.
Figure 2: Performance under different training data ratio on the LingoQA dataset. HoP surpasses the full-data performance of LLaVA-v1.5 using only 25% of the training data.
@@
Interpretability has become a crucial research focus in autonomous driving to address these concerns. Visual Question Answering (VQA) tasks have gained considerable attention as they allow systems to articulate their observations and actions through natural language interactions [41]. The application of Multimodal Large Language Models (MLLMs) [4, 51, 46, 8, 27, 14] has significantly advanced VQA tasks by integrating visual encoders with Large Language Models (LLMs) [12, 48, 10]. Typically, MLLMs comprise a visual encoder, an adapter, and an LLM. For VQA, the visual encoder extracts image features, aligned by the adapter with text tokens. The LLM tokenizes the question and integrates visual tokens to generate an answer.
VQA in autonomous driving requires capturing hierarchical visual features for complex interactions and long-tail scenarios. However, general MLLMs [4, 51, 27] struggle to capture domain-specific representations due to limited visual-language aligned data in the driving domain [43, 38]. Although existing methods [47, 42] attempt to enhance visual representation through multi-feature fusion of visual encoders like CLIP [39], DINOv2 [36], and Pix2Struct [22], their intricate fusion strategies and specialized designs fall short of addressing the unique demands of autonomous driving: domain-specific adaptability with limited data.
@@
To enable faster domain adaptation, the visual encoder must quickly capture instance-level information relevant to driving scenarios. However, as shown in Fig. 3, CLIP’s visual tokens lack token-wise affinity relationships, which are critical for preserving instance-level structure. Additionally, these general visual encoders underrepresent domain-specific semantics, making them prone to missing small, sparse, yet crucial elements for driving, such as distant vehicles, pedestrians, and traffic signs. Lastly, current MLLMs struggle to adjust visual representations based on specific questions because visual and text tokens are processed separately. This separation limits the model’s ability to focus on query-relevant image regions, reducing the effectiveness of context-aware responses in autonomous driving.
@@
Based on these observations, we introduce additional tokens as hints to enhance visual representation. Specifically, we propose three key enhancements: Affinity hint to capture instance-level structures by leveraging token-wise affinity relationships from DINOv2 [36]; Semantic hint to introduce sparse, driving-specific tokens that incorporate high-level semantic information using queries from DETR-like models; and Question hint to correlate visual features according to the query context, ensuring focus on question-relevant regions. We conduct comprehensive studies to design effective and efficient fusion strategies for combining these hint tokens. Fig. 1 illustrates our proposed method, referred to as the Hints of Prompt (HoP) framework. Fig. 2 shows that HoP outperforms the baseline with only 25% of the training data, demonstrating efficient domain adaptation with limited data in driving scenarios. Furthermore, after analyzing the gains from HoP, we propose an efficient version that directly distills hint tokens from the CLIP backbone, achieving competitive performance more efficiently.
Through extensive experiments on autonomous driving VQA benchmarks, we demonstrate that our method significantly outperforms existing approaches, achieving state-of-the-art results. Our ablation studies further confirm the effectiveness of each hints, as well as the impact of different fusion strategies and LLM scales, illustrating how each component contributes to the overall performance improvement and faster domain adaptation.
To summarize, our contributions are as follows:
@@
  • We identify key limitations of current general MLLMs with CLIP alone in autonomous driving: lack of domain priors for domain-specific visual representation.
  • We introduce the Hints of Prompt (HoP) framework, a data-efficient solution for enhancing visual representation tailored for autonomous driving scenes.
@@
  • Extensive experiments demonstrate significant improvements of our methods, establishing state-of-the-art results on LingoQA [35], DRAMA [33], and BDD-X [21].
@@ 2 Related Work
In autonomous driving, MLLMs address critical challenges in perception, planning, control, question-answering (QA), and generative world models. For example, GPT-Driver [34] reframes motion planning as a language modeling task, enhancing accuracy, while DriveGPT4 [53] combines vision and language input to predict control signals with real-time explanations, improving transparency.
In perception, MLLMs advance sensor data fusion from LiDAR and cameras; models such as Talk2BEV [11] use Bird’s Eye View (BEV) features for complex visual tasks, while LMDrive [40] and DriveVLM [45] employ trajectory-based reasoning to refine scene understanding and decision making. In QA, DriveLM [43] uses graph-based reasoning to enhance real-time planning. For world models, generative approaches like DriveDreamer [52] and GAIA-1 [19] generate realistic driving scenarios for data augmentation and training. These advancements highlight MLLMs’ potential to enhance multimodal understanding and optimize autonomous driving systems across diverse applications.
Enhancing Visual Representations in MLLMs. Effective alignment of visual representations with LLM capabilities is essential to improve overall performance. Taking advantage of data, InternVL [8] introduces a large-scale visual encoder to improve granularity of the features, and some works [31, 46, 47, 42, 26] combine multiple visual encoders with a complex fusion strategy to take advantage of complementary strengths. However, effective alignment of visual representations with LLM capabilities remains challenging in autonomous driving.
QA-ViT [17] incorporates textual information into visual features to help MLLMs focus on relevant image regions, enhancing interpretability while maintaining efficiency. Similarly, we utilize the same LLM text embeddings as Question hints, but our fusion strategy is more easily integrated with various visual backbones. While Hint-AD [15] aligns language with perception-planning tokens holistically, and TOKEN [44] leverages enriched representations generated by an end-to-end driving model to mitigate data scarcity, our method uniquely introduces dynamic hint tokens with a lightweight fusion module, enabling context-aware visual-semantic alignment for driving scenarios.
@@ 3.1 Overview
The Hints of Prompt (HoP) framework (Fig. 1) leverages three distinct token types—Affinity, Semantic, and Question hints—each representing a different level of information. These hints progress from instance-level, token-wise connections in the Affinity hint, to high-level semantic information in the Semantic hint, and finally to question-specific context in the Question hint, gradually refining and aligning the visual representation with the LLM. A simple Hint Fusion module integrates these hints into a cohesive visual representation, which is then processed by an adapter and passed to the LLM for answer generation. By incorporating hints across different information levels, our method enhances context-aware in VQA for autonomous driving and facilitates domain adaptation under data scarcity.
Figure 4: The impact of the three proposed hint types compared to the baseline and attention maps with/without these hints.
@@ 3.2 Hint Tokens
Semantic hint After implementing instance-level enhancement, providing a semantic category specific to the driving domain can further focus visual representations on key instances, such as the cyclist shown in Fig. 4. The Semantic hint aims to incorporate high-level, domain-specific semantic information, such as vehicles, traffic signs, and pedestrians, which is crucial for understanding complex interactions. To achieve this, we employ sparse queries from DETR-like [6] detection or segmentation pipelines as Semantic hints. Specifically, GroundingDINO [29] or Mask2Former [9] will be used for the extraction of Semantic hints. Experimental results (Tab. 5) show that benefiting from Mask2Former’s adaptation to driving scenarios, it can align more effectively with the task-specific requirements of VQA in autonomous driving.
@@
As shown in Fig. 4, incorporating the three proposed hints improves the model’s response over the baseline. The baseline incorrectly suggests overtaking is possible, overlooking oncoming traffic. As shown by the HoP attention map, the model with hints gains instance-level structure and focuses more accurately on key elements, such as the cyclist and oncoming vehicles. This added context enhances the model’s spatial reasoning, resulting in a more precise and relevant response in complex driving scenarios. Detailed quantitative analysis is available in Sec. 4.
Method LLM Lingo-Judge $\uparrow$ BLEU-4 $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$
GPT-4V [54] (zero-shot) - 59.6 6.30 12.4 42.8
LingoQA Baseline [35] Vicuna-v1.5-7B 60.6 15.0 18.6 65.6
QA-ViT [17] Vicuna-v1.5-7B 63.6 14.6 19.0 65.9
VTS [32] InternLM2-7B 64.2 14.5 20.5 56.9
LLaVA-v1.5 [27] Vicuna-v1.5-7B 63.2 14.1 19.3 63.7
LLaVA-v1.5 (DINOv2)† Vicuna-v1.5-7B 62.8 14.7 19.3 68.3
LLaVA-v1.5 (+A-MoF) [47] Vicuna-v1.5-7B 64.2 14.5 19.1 64.7
Efficient HoP Vicuna-v1.5-7B 66.8 15.2 20.0 66.2
HoP Vicuna-v1.5-7B 67.8 15.8 20.3 70.9
@@ 3.3 Hints Fusion
Concatenation As a baseline, hint tokens are directly concatenated with CLIP tokens, as illustrated in Fig. 5(a):
@@ 4 Experiments
Table 2: Performance comparison on the DRAMA dataset. “†” indicates models that are fine-tuned on the DRAMA dataset. HoP demonstrates enhanced performance over other methods across most metrics, with LLaVA-v1.5 as the baseline.
Method Easy Medium Hard All
@@ 4.1 Implementation Details
The whole HoP training is conducted for one epoch using the AdamW optimizer with a cosine learning rate schedule, a warm-up ratio of 0.03, and a base learning rate of 2e-5. All models are trained on 32 NVIDIA A100 GPUs with a batch size of 4 per GPU. We use an NVIDIA RTX 4090 GPU to measure latency in the inference.
@@ 4.2 Datasets and Benchmarks
LingoQA [35] is a benchmark dataset for video question answering in autonomous driving, with 28K unique driving scenarios and 419K question-answer pairs. The dataset consists of two subsets: Action and Scenery, plus a high-quality Evaluation subset for benchmarking. The main metric, Lingo-Judge, is an evaluation model aligned with human preferences, while BLEU [37], METEOR [5], and CIDEr [50] provide lexical similarity assessments.
@@
BDD-X [21] is a dataset developed for autonomous driving tasks, focusing on video captioning and rationales, we adopt the VQA pairs and evaluation protocol provided by DriveGPT4 [53], which include 26K training samples and 1.8K testing samples. BLEU, CIDEr, and ROUGE are employed as metrics.
@@
4.3 Comparison with Existing Works
Our results on the LingoQA [35], DRAMA [33], and BDD-X [21] datasets demonstrate the effectiveness of the Hints of Prompt (HoP) method in enhancing VQA for autonomous driving. As shown in Tab. 1, HoP significantly outperforms baseline on LingoQA, achieving a top Lingo-Judge score of 67.8 and leading BLEU-4, METEOR, and CIDEr scores.
As shown in Tab. 2, HoP surpasses other methods on the DRAMA dataset across key metrics, including BLEU-1, BLEU-4, METEOR, ROUGE, and SPICE, highlighting its capability to generate detailed and contextually accurate answers in urban traffic scenes.
In Tab. 3, HoP’s performance on BDD-X, evaluated using the DriveGPT4 protocol, consistently outperforms other methods across all difficulty levels (Easy, Medium, Hard, and All), achieving the highest CIDEr scores and robust results even in challenging driving scenarios.
Visualization examples can be found in Fig. 8.
@@ 4.3 Comparison with Existing Works
Table 5: Impact of different Semantic hint types. Abbreviations: SH (Semantic hint), LJ (Lingo-Judge), B4 (BLEU-4).
@@
Table 6: Impact of different Question hint types. “CLIP” refers to hint tokens derived from the CLIP text encoder, “LLM” refers to tokens from the LLM input embedding layer. Abbreviations: QH (Question Hint), LJ (Lingo-Judge), B4 (BLEU-4).